OpenStudy (anonymous):

Here is a question my formula for p=r-c is using x^2+62x - 6x+300 I had to simplify it so now it is x^2-56x+300 would that be right?

OpenStudy (anonymous):

it dont seem like it would be right jus because when i PLUG IN given numbers and in the orignal problem 300 is a cost so I would think you would have to subtract it.

OpenStudy (anonymous):

o and it is x^2+56+300 btw

OpenStudy (amistre64):

im not quite sure that you can use 'x' to indicate number of items sold AND variable costs

OpenStudy (amistre64):

now if all you are doing is buying an item and turning around and selling it; then maybe ....

OpenStudy (anonymous):

no we are plugging in numbers like 25 for example is one of the numbers but we have to simplify the

OpenStudy (anonymous):

equation

OpenStudy (anonymous):

yes buying and then selling it

OpenStudy (amistre64):

you dont have the same equations we cam up with, is this a different problem? or did you leave some things out?

OpenStudy (anonymous):

ok we we have to use the equation we come up with

OpenStudy (amistre64):

and what are they?

OpenStudy (anonymous):

c=6x+300 r= -x^2+62(x)

OpenStudy (anonymous):

P= r-c

OpenStudy (amistre64):

what is: -(6x+300) equal?

OpenStudy (anonymous):

it was a formula to solve for a input of x

OpenStudy (amistre64):

-x^2 +62x + -(6x +300) -x^2 +62x -6x -300 is what I am trying to get you to see...

OpenStudy (anonymous):

and this is waht I was thinking didnt think about -(6x+300) that would make it work lol